If 5cosec theta - 2cot theta = 5
Then find 5cot theta - 2 cosec theta

Grade:10

1 Answers

Nisarg Chadha
6 years ago
5cosecQ-2cotQ=525cosec^2Q+4cot^Q-20cosecQ.cotQ=2520cosecQ.cotQ=29cosec^Q-4-25(as cot^Q=cosec^2Q-1)So 25cot^2Q+4cosec^2Q-20cosecQcotQ=4cosec^2Q-4cot^2Q+29(using value of 20cisecQcotQ)=4+29=33